\begin{thebibliography}{100}

\bibitem{hco121}
H.~C. {\"O}ttinger.
\newblock Nonequilibrium thermodynamics -- {A} tool for applied rheologists.
\newblock {\em Appl.\ Rheol.}, 9:17--26, 1999.

\bibitem{hco206}
H.~C. {\"O}ttinger.
\newblock Nonequilibrium thermodynamics: {A} powerful tool for scientists and
  engineers.
\newblock {\em Dyna}, 79/175E:122--131, 2012.

\bibitem{hco99}
M.~Grmela and H.~C. {\"O}ttinger.
\newblock Dynamics and thermodynamics of complex fluids. {I}. {D}evelopment of
  a general formalism.
\newblock {\em Phys.\ Rev.\ E}, 56:6620--6632, 1997.

\bibitem{hco100}
H.~C. {\"O}ttinger and M.~Grmela.
\newblock Dynamics and thermodynamics of complex fluids. {II}. {I}llustrations
  of a general formalism.
\newblock {\em Phys.\ Rev.\ E}, 56:6633--6655, 1997.

\bibitem{hcobet}
H.~C. {\"O}ttinger.
\newblock {\em Beyond Equilibrium Thermodynamics}.
\newblock Wiley, Hoboken, 2005.

\bibitem{hco167}
H.~C. {\"O}ttinger.
\newblock Nonequilibrium thermodynamics of glasses.
\newblock {\em Phys.\ Rev.\ E}, 74:011113, 2006.

\bibitem{hco102}
B.~J. Edwards and H.~C. {\"O}ttinger.
\newblock Time-structure invariance criteria for closure approximations.
\newblock {\em Phys.\ Rev.\ E}, 56:4097--4103, 1997.

\bibitem{hco138}
M.~Kr{\"o}ger, M.~H{\"u}tter, and H.~C. {\"O}ttinger.
\newblock Symbolic test of the {J}acobi identity for given generalized
  ``{P}oisson'' bracket.
\newblock {\em Comput.\ Phys.\ Commun.}, 137:325--340, 2001.

\bibitem{KroegerHuetter10}
M.~Kr{\"o}ger and M.~H{\"u}tter.
\newblock Automated symbolic calculations in nonequilibrium thermodynamics.
\newblock {\em Comput.\ Phys.\ Commun.}, 181:2149--2157, 2010.

\bibitem{BerisEdwards}
A.~N. Beris and B.~J. Edwards.
\newblock {\em The Thermodynamics of Flowing Systems}.
\newblock Oxford University Press, New York, 1994.

\bibitem{Edwards98}
B.~J. Edwards.
\newblock An analysis of single and double generator thermodynamic formalisms
  for the macroscopic description of complex fluids.
\newblock {\em J.~Non-Equilib.\ Thermodyn.}, 23:301--333, 1998.

\bibitem{hco110}
B.~J. Edwards, A.~N. Beris, and H.~C. {\"O}ttinger.
\newblock An analysis of single and double generator thermodynamic formalisms
  for complex fluids. {II}. {T}he microscopic description.
\newblock {\em J.~Non-Equilib.\ Thermodyn.}, 23:334--350, 1998.

\bibitem{Beris01}
A.~N. Beris.
\newblock Bracket formulation as a source for the development of dynamic
  equations in continuum mechanics.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 96:119--136, 2001.

\bibitem{Grabert}
H.~Grabert.
\newblock {\em Projection Operator Techniques in Nonequilibrium Statistical
  Mechanics}.
\newblock Springer, Berlin, 1982.

\bibitem{Robertson66}
B.~Robertson.
\newblock Equations of motion in nonequilibrium statistical mechanics.
\newblock {\em Phys.\ Rev.}, 144:151--161, 1966.

\bibitem{Mori65}
H.~Mori.
\newblock Transport, collective motion, and {B}rownian motion.
\newblock {\em Prog.\ Theor.\ Phys.}, 33:423--455, 1965.

\bibitem{Mori65a}
H.~Mori.
\newblock A continued-fraction representation of the time-correlation
  functions.
\newblock {\em Prog.\ Theor.\ Phys.}, 34:399--416, 1965.

\bibitem{Zwanzig61}
R.~Zwanzig.
\newblock Memory effects in irreversible thermodynamics.
\newblock {\em Phys.\ Rev.}, 124:983--992, 1961.

\bibitem{hco101}
H.~C. {\"O}ttinger.
\newblock General projection operator formalism for the dynamics and
  thermodynamics of complex fluids.
\newblock {\em Phys.\ Rev.\ E}, 57:1416--1420, 1998.

\bibitem{hco131}
H.~C. {\"O}ttinger.
\newblock Derivation of two-generator framework of nonequilibrium
  thermodynamics for quantum systems.
\newblock {\em Phys.\ Rev.\ E}, 62:4720--4724, 2000.

\bibitem{hco173}
H.~C. {\"O}ttinger.
\newblock Systematic coarse graining: `{F}our lessons and a caveat' from
  nonequilibrium statistical mechanics.
\newblock {\em MRS Bulletin}, 32:936--940, 2007.

\bibitem{KuboetalII}
R.~Kubo, M.~Toda, and N.~Hashitsume.
\newblock {\em Nonequilibrium Statistical Mechanics}, volume~II of {\em
  Statistical Physics}.
\newblock Springer, Berlin, second edition, 1991.

\bibitem{hco117}
H.~C. {\"O}ttinger and A.~N. Beris.
\newblock Thermodynamically consistent reptation model without independent
  alignment.
\newblock {\em J.~Chem.\ Phys.}, 110:6593--6596, 1999.

\bibitem{hco119}
H.~C. {\"O}ttinger.
\newblock Thermodynamically admissible reptation models with anisotropic tube
  cross sections and convective constraint release.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 89:165--185, 2000.

\bibitem{hco123}
H.~C. {\"O}ttinger.
\newblock A thermodynamically admissible reptation model for fast flows of
  entangled polymers.
\newblock {\em J.~Rheol.}, 43:1461--1493, 1999.

\bibitem{hco129}
J.~Fang, M.~Kr{\"o}ger, and H.~C. {\"O}ttinger.
\newblock A thermodynamically admissible reptation model for fast flows of
  entangled polymers. {II}. {M}odel predictions for shear and elongational
  flows.
\newblock {\em J.~Rheol.}, 44:1293--1317, 2000.

\bibitem{EslamiGrmela08}
H.~Eslami and M.~Grmela.
\newblock Mesoscopic formulation of reptation.
\newblock {\em Rheol.\ Acta}, 47:399--415, 2008.

\bibitem{hco210}
T.~Savin, W.~J. Briels, and H.~C. {\"O}ttinger.
\newblock Thermodynamic formulation of flowing soft matter with transient
  forces.
\newblock {\em Rheol.\ Acta}, 52:23--32, 2013.

\bibitem{Stephanouetal16}
P.~S. Stephanou, I.~C. Tsimouri, and V.~G. Mavrantzas.
\newblock Flow-induced orientation and stretching of entangled polymers in the
  framework of nonequilibrium thermodynamics.
\newblock {\em Macromolecules}, 49:3161--3173, 2016.

\bibitem{hco130}
H.~C. {\"O}ttinger.
\newblock Thermodynamic admissibility of the pompon model for branched
  polymers.
\newblock {\em Rheol.\ Acta}, 40:317--321, 2001.

\bibitem{vanMeerveld02}
J.~{van Meerveld}.
\newblock Note on thermodynamic consistency of the integral pom-pom model.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 108:291--299, 2002.

\bibitem{hco108}
N.~J. Wagner, H.~C. {\"O}ttinger, and B.~J. Edwards.
\newblock Generalized {D}oi-{O}hta model for multiphase flow developed via
  {GENERIC}.
\newblock {\em AIChE J.}, 45:1169--1181, 1999.

\bibitem{GrmelaBousPal01}
M.~Grmela, M.~Bousmina, and J.-F. Palierne.
\newblock On the rheology of immiscible blends.
\newblock {\em Rheol.\ Acta}, 40:560--569, 2001.

\bibitem{DresslerEdwards04}
M.~Dressler and B.~J. Edwards.
\newblock The influence of matrix viscoelasticity on the rheology of polymer
  blends.
\newblock {\em Rheol.\ Acta}, 43:257--282, 2004.

\bibitem{GuGrmela08}
J.~F. Gu and M.~Grmela.
\newblock Flow properties of immiscible blends: {D}oi-{O}hta model with active
  advection.
\newblock {\em Phys.\ Rev.\ E}, 78:056302, 1--17, 2008.

\bibitem{GuGrmelaBous08}
J.~F. Gu, M.~Grmela, and M.~Bousmina.
\newblock A mesoscopic rheological model of immiscible blends with the
  interface covered with a surface active agent.
\newblock {\em Phys.\ Fluids}, 20:043102, 1--20, 2008.

\bibitem{hco193}
A.~Jeli{\'c}, P.~Ilg, and H.~C. {\"O}ttinger.
\newblock Bridging length and time scales in sheared demixing systems: {F}rom
  the {C}ahn-{H}illiard to the {D}oi-{O}hta model.
\newblock {\em Phys.\ Rev.\ E}, 81:011131, 2010.

\bibitem{WagnerN01}
N.~J. Wagner.
\newblock The {S}moluchowski equation for colloidal suspensions developed and
  analyzed through the {GENERIC} formalism.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 96:177--201, 2001.

\bibitem{GrmelaAitKadiLafl98}
M.~Grmela, A.~A{\"\i}t-Kadi, and P.~G. Lafleur.
\newblock Suspensions of fibres in viscoelastic fluids: Rheology.
\newblock {\em J.~Chem.\ Phys.}, 109:6973--6981, 1998.

\bibitem{Keshtkaretal10}
M.~Keshtkar, M.-C. Heuzey, P.~J. Carreau, M.~Rajabian, and C.~Dubois.
\newblock Rheological properties and microstructural evolution of semi-flexible
  fiber suspensions under shear flow.
\newblock {\em J.~Rheol.}, 54:197--222, 2010.

\bibitem{Grmelaetal14}
M.~Grmela, A.~Ammar, F.~Chinesta, and G.~Maitrejean.
\newblock A mesoscopic rheological model of moderately concentrated colloids.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 212:1--12, 2014.

\bibitem{Hutter01}
M.~H{\"u}tter.
\newblock Thermodynamically consistent incorporation of the {S}chneider rate
  equations into two-phase models.
\newblock {\em Phys.\ Rev.\ E}, 64:011209, 1--11, 2001.

\bibitem{Hutter02}
M.~H{\"u}tter.
\newblock {GENERIC} treatment of compressible two-phase flow: Convection
  mechanism of scalar morphological variables.
\newblock {\em J.~Non-Equilib.\ Thermodyn.}, 27:349--365, 2002.

\bibitem{Espanol01}
P.~Espa{\~n}ol.
\newblock Thermodynamics for a van der {W}aals fluid.
\newblock {\em J.~Chem.\ Phys.}, 115:5392--5403, 2001.

\bibitem{EspanolThieu03}
P.~Espa{\~n}ol and C.~Thieulot.
\newblock Microscopic derivation of hydrodynamic equations for phase-separating
  fluid mixtures.
\newblock {\em J.~Chem.\ Phys.}, 118:9109--9127, 2003.

\bibitem{hco109}
H.~C. {\"O}ttinger.
\newblock On the structural compatibility of a general formalism for
  nonequilibrium dynamics with special relativity.
\newblock {\em Physica A}, 259:24--42, 1998.

\bibitem{hco111}
H.~C. {\"O}ttinger.
\newblock Relativistic and nonrelativistic description of fluids with
  anisotropic heat conduction.
\newblock {\em Physica A}, 254:433--450, 1998.

\bibitem{hco112}
H.~C. {\"O}ttinger.
\newblock Thermodynamically admissible equations for causal dissipative
  cosmology, galaxy formation, and transport processes in a gravitational
  collapse.
\newblock {\em Phys.\ Rev.\ D}, 60:103507, 1999.

\bibitem{hco114}
P.~Ilg and H.~C. {\"O}ttinger.
\newblock Nonequilibrium relativistic thermodynamics in bulk viscous cosmology.
\newblock {\em Phys.\ Rev.\ D}, 61:023510, 1999.

\bibitem{hco180}
H.~C. {\"O}ttinger.
\newblock Role of nonequilibrium entropy in {E}instein's theory of gravitation.
\newblock {\em Physica A}, 387:4560--4564, 2008.

\bibitem{Duong15}
M.~H. Duong.
\newblock Formulation of the relativistic heat equation and the relativistic
  kinetic {F}okker-{P}lanck equations using {GENERIC}.
\newblock {\em Physica A}, 439:34--43, 2015.

\bibitem{hco122}
P.~Espa{\~n}ol, M.~Serrano, and H.~C. {\"O}ttinger.
\newblock Thermodynamically admissible form for discrete hydrodynamics.
\newblock {\em Phys.\ Rev.\ Lett.}, 83:4542--4545, 1999.

\bibitem{SerranoEspanol01}
M.~Serrano and P.~Espa{\~n}ol.
\newblock Thermodynamically consistent mesoscopic fluid particle model.
\newblock {\em Phys.\ Rev.\ E}, 64:046115, 1--18, 2001.

\bibitem{EspanolRev03}
P.~Espa{\~n}ol and M.~Revenga.
\newblock Smoothed dissipative particle dynamics.
\newblock {\em Phys.\ Rev.\ E}, 67:026705, 1--12, 2003.

\bibitem{ElleroEspanolFlek03}
M.~Ellero, P.~Espa{\~n}ol, and E.~G. Flekk{\o}y.
\newblock Thermodynamically consistent fluid particle model for viscoelastic
  flows.
\newblock {\em Phys.\ Rev.\ E}, 68:041504, 1--19, 2003.

\bibitem{hco142}
M.~Kr{\"o}ger, J.~Ram{\'\i}rez, and H.~C. {\"O}ttinger.
\newblock Projection from an atomistic chain contour to its primitive path.
\newblock {\em Polymer}, 43:477--487, 2002.

\bibitem{hco146}
V.~G. Mavrantzas and H.~C. {\"O}ttinger.
\newblock Atomistic {M}onte {C}arlo simulations of polymer melt elasticity:
  Their nonequilibrium thermodynamics {GENERIC} formulation in a generalized
  canonical ensemble.
\newblock {\em Macromolecules}, 35:960--975, 2002.

\bibitem{hco145}
V.~A. Harmandaris, V.~G. Mavrantzas, D.~N. Theodorou, M.~Kr{\"o}ger,
  J.~Ram{\'\i}rez, H.~C. {\"O}ttinger, and D.~Vlassopoulos.
\newblock Crossover from the {R}ouse to the entangled polymer melt regime:
  Signals from long, detailed atomistic molecular dynamics simulations,
  supported by rheological experiments.
\newblock {\em Macromolecules}, 36:1376--1387, 2003.

\bibitem{Kroeger04}
M.~Kr{\"o}ger.
\newblock Simple models for complex nonequilibrium fluids.
\newblock {\em Phys.\ Rep.}, 390:453--551, 2004.

\bibitem{hco155}
H.~C. {\"O}ttinger.
\newblock Coarse-graining of wormlike polymer chains for substantiating
  reptation.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 120:207--213, 2004.

\bibitem{hco157}
M.~Kr{\"o}ger and H.~C. {\"O}ttinger.
\newblock Beyond-equilibrium molecular dynamics of a rarefied gas subjected to
  shear flow.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 120:175--187, 2004.

\bibitem{hco182}
P.~Ilg, H.~C. {\"O}ttinger, and M.~Kr{\"o}ger.
\newblock Systematic time-scale-bridging molecular dynamics applied to flowing
  polymer melts.
\newblock {\em Phys.\ Rev.\ E}, 79:011802, 2009.

\bibitem{IlgKroger11}
P.~Ilg and M.~Kr{\"o}ger.
\newblock Molecularly derived constitutive equation for low-molecular polymer
  melts from thermodynamically guided simulation.
\newblock {\em J.~Rheol.}, 55:69--93, 2011.

\bibitem{Ilgetal11}
P.~Ilg, M.~H{\"u}tter, and M.~Kr{\"o}ger.
\newblock Ideal contribution to the macroscopic, quasiequilibrium entropy of
  anisotropic fluids.
\newblock {\em Phys.\ Rev.\ E}, 83:061713, 1--7, 2011.

\bibitem{hco195}
C.~Baig, V.~G. Mavrantzas, and H.~C. {\"O}ttinger.
\newblock On {M}axwell’s relations of thermodynamics for polymeric liquids away
  from equilibrium.
\newblock {\em Macromolecules}, 44:640--646, 2011.

\bibitem{LiuDeKee05a}
Q.~Liu and D.~{De Kee}.
\newblock Modeling of diffusion through polymeric membranes.
\newblock {\em Rheol.\ Acta}, 44:287--294, 2005.

\bibitem{LiuDeKee05b}
Q.~Liu and D.~{De Kee}.
\newblock Modeling of diffusion through nanocomposite membranes.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 131:32--43, 2005.

\bibitem{GrmelaLebDub11}
M.~Grmela, G.~Lebon, and C.~Dubois.
\newblock Multiscale thermodynamics and mechanics of heat.
\newblock {\em Phys.\ Rev.\ E}, 83:061134, 1--15, 2011.

\bibitem{GuGrmelaBous10}
J.~F. Gu, M.~Grmela, and M.~Bousmina.
\newblock Mesohydrodynamics of membrane suspensions.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 165:75--83, 2010.

\bibitem{hco194}
H.~C. {\"O}ttinger.
\newblock Thermodynamically admissible 13 moment equations from the {B}oltzmann
  equation.
\newblock {\em Phys.\ Rev.\ Lett.}, 104:120601, 2010.

\bibitem{StruchtrupTorr10}
H.~Struchtrup and M.~Torrilhon.
\newblock Comment on `thermodynamically admissible 13 moment equations from the
  {B}oltzmann equation'.
\newblock {\em Phys.\ Rev.\ Lett.}, 105:128901, 2010.

\bibitem{hco198}
H.~C. {\"O}ttinger.
\newblock {\"O}ttinger replies.
\newblock {\em Phys.\ Rev.\ Lett.}, 105:128902, 2010.

\bibitem{hco216}
S.~Kjelstrup, D.~Bedeaux, and H.~C. {\"O}ttinger.
\newblock Nonlinear coupled equations for electrochemical cells as developed by
  the general equation for nonequilibrium reversible-irreversible coupling.
\newblock {\em J.~Chem.\ Phys.}, 141:124102, 2014.

\bibitem{Mukherjeeetal04}
J.~Mukherjee, S.~Wilson, and A.~N. Beris.
\newblock Flow-induced nonequilibrium thermodynamics of lamellar
  semicrystalline polymers.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 120:225--240, 2004.

\bibitem{Hutter04}
M.~H{\"u}tter.
\newblock Crystallization under external pressure.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 120:55--68, 2004.

\bibitem{HutterRutArm05}
M.~H{\"u}tter, G.~C. Rutledge, and R.~C. Armstrong.
\newblock Crystal shapes and crystallization in continuum modeling.
\newblock {\em Phys.\ Fluids}, 17:014107, 1--13, 2005.

\bibitem{Hutter06}
M.~H{\"u}tter.
\newblock Volume change and non-local driving force in crystallization.
\newblock {\em J.~Non-Equilib.\ Thermodyn.}, 31:73--101, 2006.

\bibitem{vanMeerveldetal08}
J.~{van Meerveld}, M.~H{\"u}tter, and G.~W.~M. Peters.
\newblock Continuum model for the simulation of fiber spinning, with quiescent
  and flow-induced crystallization.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 150:177--195, 2008.

\bibitem{MukherjeeBeris10}
J.~Mukherjee and A.~N. Beris.
\newblock Multiscale modeling of crystallization morphologies in high speed
  fiber spinning of semicrystalline polymers.
\newblock {\em J.~Comput.\ Theor.\ Nanosci.}, 7:726--737, 2010.

\bibitem{BaigEdwards10a}
C.~Baig and B.~J. Edwards.
\newblock Atomistic simulation of flow-induced crystallization at constant
  temperature.
\newblock {\em Europhys.\ Lett.}, 89:36003, 1--6, 2010.

\bibitem{BaigEdwards10b}
C.~Baig and B.~J. Edwards.
\newblock Atomistic simulation of crystallization of a polyethylene melt in
  steady uniaxial extension.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 165:992--1004, 2010.

\bibitem{hco176}
M.~H{\"u}tter, M.~Grmela, and H.~C. {\"O}ttinger.
\newblock What is behind the plastic strain rate?
\newblock {\em Rheol.\ Acta}, 48:769--778, 2009.

\bibitem{HutterTervoort08b}
M.~H{\"u}tter and T.~A. Tervoort.
\newblock Finite anisotropic elasticity and material frame indifference from a
  nonequilibrium thermodynamics perspective.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 152:45--52, 2008.

\bibitem{HutterTervoort08}
M.~H{\"u}tter and T.~A. Tervoort.
\newblock Thermodynamic considerations on non-isothermal finite anisotropic
  elasto-viscoplasticity.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 152:53--65, 2008.

\bibitem{hco192}
H.~C. {\"O}ttinger.
\newblock {L}ie groups in nonequilibrium thermodynamics: {G}eometric structure
  behind viscoplasticity.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 165:1005--1013, 2010.

\bibitem{Mielke11}
A.~Mielke.
\newblock Formulation of thermoelastic dissipative material behavior using
  {GENERIC}.
\newblock {\em Continuum Mech.\ Thermodyn.}, 23:233--256, 2011.

\bibitem{HuttervBremen12}
M.~H{\"u}tter and L.~C.~A. {van Bremen}.
\newblock Microstructural model for the plasticity of amorphous solids.
\newblock {\em J.~Appl.\ Polym.\ Sci.}, 125:4376--4389, 2012.

\bibitem{HutterSvendsen12}
M.~H{\"u}tter and B.~Svendsen.
\newblock Thermodynamic model formulation for viscoplastic solids as general
  equations for non-equilibrium reversible-irreversible coupling.
\newblock {\em Continuum Mech.\ Thermodyn.}, 24:211--227, 2012.

\bibitem{HutterTervoort15}
M.~H{\"u}tter and T.~A. Tervoort.
\newblock Statistical-mechanics based modeling of anisotropic viscoplastic
  deformation.
\newblock {\em Mech.\ Mater.}, 80:37--51, 2015.

\bibitem{Peshkovetal15}
I.~Peshkov, M.~Grmela, and E.~Romenski.
\newblock Irreversible mechanics and thermodynamics of two-phase continua
  experiencing stress-induced solid-fluid transitions.
\newblock {\em Continuum Mech.\ Thermodyn.}, 27:905--940, 2015.

\bibitem{HutterTervoort08a}
M.~H{\"u}tter and T.~A. Tervoort.
\newblock Continuum damage mechanics: {C}ombining thermodynamics with a
  thoughtful characterization of the microstructure.
\newblock {\em Acta Mech.}, 201:297--312, 2008.

\bibitem{hco177}
E.~{Del Gado}, P.~Ilg, M.~Kr{\"o}ger, and H.~C. {\"O}ttinger.
\newblock Structure as a static signature of cooperativity in supercooled
  liquids.
\newblock {\em Phys.\ Rev.\ Lett.}, 101:095501, 2008.

\bibitem{hco196}
M.~Mosayebi, E.~{Del Gado}, P.~Ilg, and H.~C. {\"O}ttinger.
\newblock Probing a critical length scale at the glass transition.
\newblock {\em Phys.\ Rev.\ Lett.}, 104:205704, 2010.

\bibitem{hco208}
M.~Mosayebi, E.~{Del Gado}, P.~Ilg, and H.~C. {\"O}ttinger.
\newblock Deformation of inherent structures to detect long-range correlations
  in supercooled liquids.
\newblock {\em J.~Chem.\ Phys.}, 137:024504, 2012.

\bibitem{hco213}
A.~M. Luo and H.~C. {\"O}ttinger.
\newblock Nonequilibrium thermodynamics of the shear-transformation-zone model.
\newblock {\em Phys.\ Rev.\ E}, 89:022137, 2014.

\bibitem{FuerederIlg13}
I.~Fuereder and P.~Ilg.
\newblock Nonequilibrium thermodynamics of the soft glassy rheology model.
\newblock {\em Phys.\ Rev.\ E}, 88:042134, 2013.

\bibitem{SemkivHutter16}
M.~Semkiv and M.~H{\"u}tter.
\newblock Modeling aging and mechanical rejuvenation of amorphous solids.
\newblock {\em J.~Non-Equilib.\ Thermodyn.}, 41:79--88, 2016.

\bibitem{Waldmann67}
L.~Waldmann.
\newblock Non-equilibrium thermodynamics of boundary conditions.
\newblock {\em Z.~Naturforsch.~A}, 22:1269--1286, 1967.

\bibitem{BrennerGane00}
H.~Brenner and V.~Ganesan.
\newblock Molecular wall effects: Are conditions at a boundary ``boundary
  conditions''?
\newblock {\em Phys.\ Rev.\ E}, 61:6879--6897, 2000.

\bibitem{hco162}
H.~C. {\"O}ttinger.
\newblock Nonequilibrium thermodynamics for open systems.
\newblock {\em Phys.\ Rev.\ E}, 73:036126, 2006.

\bibitem{hco171}
A.~N. Beris and H.~C. {\"O}ttinger.
\newblock Bracket formulation of nonequilibrium thermodynamics for systems
  interacting with the environment.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 152:2--11, 2008.

\bibitem{hco172}
H.~C. {\"O}ttinger.
\newblock Thermodynamic formulation of wall slip.
\newblock {\em J.~Non-Newtonian Fluid Mech.}, 152:66--75, 2008.

\bibitem{BedeauxAlbMazur76}
D.~Bedeaux, A.~M. Albano, and P.~Mazur.
\newblock Boundary conditions and non-equilibrium thermodynamics.
\newblock {\em Physica A}, 82:438--462, 1976.

\bibitem{Bedeaux86}
D.~Bedeaux.
\newblock Nonequilibrium thermodynamics and statistical physics of surfaces.
\newblock {\em Adv.\ Chem.\ Phys.}, 64:47--109, 1986.

\bibitem{AlbanoBedeaux87}
A.~M. Albano and D.~Bedeaux.
\newblock Non-equilibrium electro-thermodynamics of polarizable multicomponent
  fluids with an interface.
\newblock {\em Physica A}, 147:407--435, 1987.

\bibitem{JohannessenBed03}
E.~Johannessen and D.~Bedeaux.
\newblock The nonequilibrium van der {W}aals square gradient model. ({II}).
  {L}ocal equilibrium of the {G}ibbs surface.
\newblock {\em Physica A}, 330:354--372, 2003.

\bibitem{JohannessenGroBed08}
E.~Johannessen, J.~Gross, and D.~Bedeaux.
\newblock Nonequilibrium thermodynamics of interfaces using classical density
  functional theory.
\newblock {\em J.~Chem.\ Phys.}, 129:184703, 1--14, 2008.

\bibitem{GlavatskiyBed09}
K.~S. Glavatskiy and D.~Bedeaux.
\newblock Numerical solution of the non-equilibrium square gradient model and
  verification of local equilibrium for the {G}ibbs surface in a two-phase
  binary mixture.
\newblock {\em Phys.\ Rev.\ E}, 79:031608, 1--19, 2009.

\bibitem{hco204}
T.~Savin, K.~S. Glavatskiy, S.~Kjelstrup, H.~C. {\"O}ttinger, and D.~Bedeaux.
\newblock Local equilibrium of the {G}ibbs interface in two-phase systems.
\newblock {\em Europhys.\ Lett.}, 97:40002, 2012.

\bibitem{hco219}
M.~Schweizer, H.~C. {\"O}ttinger, and T.~Savin.
\newblock Nonequilibrium thermodynamics of an interface.
\newblock {\em Phys.\ Rev.\ E}, 93:052803, 2016.

\bibitem{hco188}
H.~C. {\"O}ttinger, D.~Bedeaux, and D.~C. Venerus.
\newblock Nonequilibrium thermodynamics of transport through moving interfaces
  with application to bubble growth and collapse.
\newblock {\em Phys.\ Rev.\ E}, 80:021606, 2009.

\bibitem{Sagis10}
L.~M.~C. Sagis.
\newblock {GENERIC} model for multiphase systems.
\newblock {\em Adv.\ Colloid Interface Sci.}, 153:58--69, 2010.

\bibitem{Sagis10a}
L.~M.~C. Sagis.
\newblock Nonlinear rheological models for structured interfaces.
\newblock {\em Physica A}, 389:1993--2006, 2010.

\bibitem{Sagis11}
L.~M.~C. Sagis.
\newblock Dynamic properties of interfaces in soft matter: {E}xperiments and
  theory.
\newblock {\em Rev.\ Mod.\ Phys.}, 83:1367--1403, 2011.

\bibitem{hco207}
L.~M.~C. Sagis and H.~C. {\"O}ttinger.
\newblock Dynamics of multiphase systems with complex microstructure. {I}.
  {D}evelopment of the governing equations through nonequilibrium
  thermodynamics.
\newblock {\em Phys.\ Rev.\ E}, 88:022149, 2013.

\bibitem{Sagis13}
L.~M.~C. Sagis.
\newblock Dynamics of multiphase systems with complex microstructure. {II}.
  {P}article-stabilized interfaces.
\newblock {\em Phys.\ Rev.\ E}, 88:022150, 2013.

\bibitem{Sagis14}
L.~M.~C. Sagis.
\newblock Dynamic behavior of interfaces: {M}odeling with nonequilibrium
  thermodynamics.
\newblock {\em Adv.\ Colloid Interface Sci.}, 206:328--343, 2014.

\bibitem{SchweizerSagis14}
M.~Schweizer and L.~M.~C. Sagis.
\newblock Nonequilibrium thermodynamics of nucleation.
\newblock {\em J.~Chem.\ Phys.}, 141:224102, 2014.

\bibitem{hco223}
A.~M. Luo, L.~M.~C. Sagis, H.~C. {\"O}ttinger, C.~{De Michele}, and P.~Ilg.
\newblock Modelling the rheology of anisotropic particles adsorbed on a
  two-dimensional fluid interface.
\newblock {\em Soft Matter}, 11:4383--4395, 2015.

\bibitem{BreuerPetru}
H.-P. Breuer and F.~Petruccione.
\newblock {\em The Theory of Open Quantum Systems}.
\newblock Oxford University Press, Oxford, 2002.

\bibitem{Weiss}
U.~Weiss.
\newblock {\em Quantum Dissipative Systems}.
\newblock Series in Modern Condensed Matter Physics, Volume~13. World
  Scientific, Singapore, third edition, 2008.

\bibitem{Grabert82}
H.~Grabert.
\newblock Nonlinear relaxation and fluctuations of damped quantum systems.
\newblock {\em Z.~Physik B}, 49:161--172, 1982.

\bibitem{hco199}
H.~C. {\"O}ttinger.
\newblock The geometry and thermodynamics of dissipative quantum systems.
\newblock {\em Europhys.\ Lett.}, 94:10006, 2011.

\bibitem{hco221}
D.~Taj and H.~C. {\"O}ttinger.
\newblock Natural approach to quantum dissipation.
\newblock {\em Phys.\ Rev.\ A}, 92:062128, 2015.

\bibitem{hco201}
H.~C. {\"O}ttinger.
\newblock Nonlinear thermodynamic quantum master equation: {P}roperties and
  examples.
\newblock {\em Phys.\ Rev.\ A}, 82:052119, 2010.

\bibitem{hco200}
H.~C. {\"O}ttinger.
\newblock Dynamic coarse-graining approach to quantum field theory.
\newblock {\em Phys.\ Rev.\ D}, 84:065007, 2011.

\bibitem{hco209}
M.~Osmanov and H.~C. {\"O}ttinger.
\newblock Open quantum systems coupled to time-dependent classical
  environments.
\newblock {\em Int.\ J.\ of Thermodynamics}, 34:1255--1264, 2013.

\bibitem{hco217}
J.~Flakowski, M.~Osmanov, D.~Taj, and H.~C. {\"O}ttinger.
\newblock Time-driven quantum master equations and their compatibility with the
  fluctuation dissipation theorem.
\newblock {\em Phys.\ Rev.\ A}, 90:042110, 2014.

\bibitem{hco222}
J.~Flakowski, M.~Osmanov, D.~Taj, and H.~C. {\"O}ttinger.
\newblock Biexponential decay and ultralong coherence of a qubit.
\newblock {\em Europhys.\ Lett.}, 113:40003, 2016.

\bibitem{hco202}
H.~C. {\"O}ttinger.
\newblock Stochastic process behind nonlinear thermodynamic quantum master
  equation. {I}. {M}ean-field construction.
\newblock {\em Phys.\ Rev.\ A}, 86:032101, 2012.

\bibitem{hco205}
J.~Flakowski, M.~Schweizer, and H.~C. {\"O}ttinger.
\newblock Stochastic process behind nonlinear thermodynamic quantum master
  equation. {II}. {S}imulation.
\newblock {\em Phys.\ Rev.\ A}, 86:032102, 2012.

\bibitem{SanzSernaCalvo}
J.~M. Sanz-Serna and M.~P. Calvo.
\newblock {\em Numerical Hamiltonian Problems}.
\newblock Applied Mathematics and Mathematical Computation, Volume~7. Chapman
  \&\ Hall, London, 1994.

\bibitem{Moser68}
J.~Moser.
\newblock Lectures on {H}amiltonian systems.
\newblock {\em Mem.\ Am.\ Math.\ Soc.}, 81:1--60, 1968.

\end{thebibliography}

\documentclass[]{jetc2017abstract}
\usepackage{graphicx}
\usepackage{times}

\begin{document}

\vbox{\hbox to \textwidth{\begin{tabular}{@{}c@{\hskip3pc}}
        \hbox to 46pt{\vbox to 3pt{\vss\hsize46pt\vss}}\end{tabular}\hss
        \vbox{\hsize37pc\scriptsize\sf\vskip\baselineskip%
            \hfill \textit {14th Joint European Thermodynamics Conference}\par
            \hfill \textit{Budapest, May 21--25, 2017}\par}
        \hskip1pc}
        }
\vskip15mm

\setcounter{page}{73}

\begin{center}
\textbf{\large\uppercase{GENERIC: Review of successful applications and a challenge for the future}}

\addvspace{12pt}\noindent

\underline{Hans Christian \"Ottinger}$^{1,*}$

\addvspace{10pt}\noindent

$^1$ETH Z\"urich, Department of Materials, Polymer Physics, CH-8093 Z\"urich, Switzerland\\
$^*$hco@mat.ethz.ch\\

\end{center}
%\addvspace{20pt}\noindent
\addvspace{4mm}\noindent

\small

\section*{ABSTRACT}

We present and discuss the two-generator framework of nonequilibrium thermodynamics known as GENERIC (``general equation for the nonequilibrium reversible-irreversible coupling''), which is based on geometric concepts and on statistical mechanics. The usefulness and maturity of the framework are illustrated by reviewing a large number of successful applications. Finally, we offer an important challenge for the future.

This abstract is based on the compact reviews \cite{hco121,hco206} presenting modern nonequilibrium thermodynamics to applied scientists and engineers. In the spirit of a continuation of these reviews, we here add some new developments of the last five years.

\subsection*{GENERIC framework}
Time-evolution equations for nonequilibrium systems possess a well-defined thermodynamic structure in which reversible and irreversible contributions are constructed separately. The reversible contribution is assumed to be of the Hamiltonian form (driven by the gradient of energy) and hence requires an underlying geometric structure which reflects the idea that the reversible time evolution should be ``under mechanistic control.'' The remaining irreversible contribution is assumed to be driven by the gradient of a nonequilibrium entropy. Our discussion is based on the GENERIC framework for closed nonequilibrium systems \cite{hco99,hco100,hcobet} (see also the concise summary in \cite{hco167} adapted for our purposes here),
\begin{equation} \label{LMformulation}
  \frac{dx}{dt} = L \cdot \frac{\delta E}{\delta x} +
  M \cdot \frac{\delta S}{\delta x} ,
\end{equation}
where $x$ represents the set of independent variables required for a complete description of a given closed nonequilibrium system, $E$ and $S$ are the total energy and entropy expressed in terms of the system variables $x$, and $L$ and $M$ are certain linear operators, or matrices. The so-called Poisson matrix $L$ and the friction matrix $M$ can also depend on $x$ so that the fundamental evolution equation (\ref{LMformulation}) can be highly nonlinear. The two contributions to the time evolution of $x$ generated by the total energy $E$ and the entropy $S$ in (\ref{LMformulation}) are the reversible and irreversible contributions to dynamics, respectively. Because $x$ typically contains position-dependent fields, such as the local mass, momentum and energy densities of hydrodynamics, the state variables are usually labeled by continuous (position) labels in addition to discrete ones. A matrix multiplication, which can alternatively be considered as the application of a linear operator, hence implies not only summations over discrete indices but also integrations over continuous labels, and the gradient $\delta/\delta x$ typically implies functional rather than partial derivatives. Equation (\ref{LMformulation}) is supplemented by the complementary degeneracy requirements
\begin{equation} \label{LSconsistency}
  L \cdot \frac{\delta S}{\delta x}=0
\end{equation}
and
\begin{equation} \label{MEconsistency}
  M \cdot \frac{\delta E}{\delta x}=0 .
\end{equation}
The requirement that the entropy gradient $\delta S/\delta x$ is in the null-space of the Poisson matrix $L$ in (\ref{LSconsistency}) expresses the reversible nature of the first contribution to the dynamics, irrespective of the particular form of the Hamiltonian. The requirement that the energy gradient $\delta E/\delta x$ is in the null-space of the friction matrix $M$ in (\ref{MEconsistency}) expresses the conservation of the total energy in a closed system by the irreversible contribution to the dynamics. Furthermore, it is required that the matrix $L$ is antisymmetric, whereas $M$ is Onsager-Casimir symmetric (see Section 3.2 of \cite{hcobet} for details) and positive-semidefinite. Positive-semidefiniteness implies the second law of nonequilibrium thermodynamics,
\begin{equation} \label{net2ndlaw}
  \frac{dS}{dt} = \frac{\delta S}{\delta x} \cdot
  M \cdot \frac{\delta S}{\delta x} \geq 0 ,
\end{equation}
which follows directly from the chain rule by using (\ref{LSconsistency}) and (\ref{MEconsistency}). Finally, the Poisson bracket that can be associated with the antisymmetric matrix $L$ is assumed to satisfy the Jacobi identity, which expresses the time-structure invariance of the reversible dynamics \cite{hco102} and can be conveniently and rigorously tested by using symbolic mathematical tools \cite{hco138,KroegerHuetter10}.

Note that energy and entropy are the fundamental concepts of nonequilibrium thermodynamics. In cases where a nonequilibrium temperature can be introduced in a meaningful way, at least locally, one may be able to combine the two generators $E$ and $S$ into a single one (a ``nonequilibrium free energy''). The possibilities and limitations of the single-generator approach have been discussed in the textbook \cite{BerisEdwards} and in detailed comparisons to the double-generator approach \cite{Edwards98,hco110,Beris01}.

A very reassuring feature of the GENERIC framework is that there exists a clear link to nonequilibrium statistical mechanics. By eliminating the fast degrees of freedom, the projection-operator formalism \cite{Grabert,Robertson66,Mori65,Mori65a,Zwanzig61} produces equations of the GENERIC form (see Chapter~6 of \cite{hcobet} and \cite{hco101,hco131,hco173}). The fast degrees result in noise and friction felt by the slow variables, where these two effects are found to be intimately related according to the fluctuation-dissipation theorem (see Section~1.6 of \cite{KuboetalII}). As a result of the projection procedure, well-defined statistical expressions for the thermodynamic building blocks $E,S,L,M$ in (\ref{LMformulation})-(\ref{MEconsistency}) arise (see Section~6.1.4 of \cite{hcobet}). Evaluation of these expressions should be the Holy Grail of computer simulations for nonequilibrium systems \cite{hco173}. The counterpart in equilibrium statistical thermodynamics is the determination of partition functions (or their partial derivatives) by Monte Carlo simulations to obtain thermodynamic information in terms of the free energy.

A cornerstone of nonequilibrium statistical mechanics is the nonequilibrium ensemble. It is a probability density on the larger space of more microscopic states, parametrized by the more macroscopic state variables taking values from a smaller space (to avoid awkward formulations, we simply refer to microscopic and macroscopic states from now on). The ideas of microcanonical, canonical, and mixed ensembles are taken over from equilibrium statistical mechanics, but now with a much larger and less universal set of thermodynamic nonequilibrium variables. The energy $E$ and the Poisson matrix $L$ of the coarse-grained description can be obtained by simply averaging their miscroscopic counterparts by means of the nonequilibrium ensemble. As at equilibrium, the evaluation of the entropy $S$ is a matter of counting microscopic states or, more generally, of properly normalizing the ensemble. The friction matrix $M$ is the only building block that requires dynamic material information. According to the fluctuation-dissipation theorem, it can be obtained from the time-correlation functions of fluctuations. More precisely, one needs to evaluate the time-integral of two-time correlations of the fluctuations of the macroscopic variables resulting from the elimination of fast microscopic degrees of freedom. The explicit expression for the friction matrix is known as the Green-Kubo formula (see, for example, (3.47) or (6.73) of \cite{hcobet} or (3) of \cite{hco173}).

The respective role and potential of Monte Carlo, molecular dynamics, and Brownian dynamics simulations in thermodynamically guided simulations for nonequilibrium systems has been elaborated in \cite{hco173}. Dynamic simulations should run only over a fraction of the characteristic slow time scales, just sufficiently long to evaluate the decay of two-time correlations on the fast time scales. Initial conditions should be obtained by Monte Carlo sampling from nonequilibrium ensembles. Thermostats and similar devices are unnecessary for such short simulations.

\subsection*{Review of applications}
In addition to the many famous applications of linear irreversible thermodynamics (such as diffusion, osmotic pressure, heat conduction, propagation of sound, electrokinetic effects, thermoelectric effects, thermokinetic effects, dielectric relaxation, polarizable media in electromagnetic fields, magneto-plasmas, superfluids, viscoelastic fluids, and chemical reactions), a number of applications of nonlinear irreversible thermodynamics have been compiled in Appendix~E of the textbook \cite{hcobet}. Those advanced applications are predominantly from the field of complex fluids: reptation model for entangled linear polymers \cite{hco117,hco119,hco123,hco129,EslamiGrmela08,hco210,Stephanouetal16}, pompon model for branched polymers \cite{hco130,vanMeerveld02}, polymer blends \cite{hco108,GrmelaBousPal01,DresslerEdwards04,GuGrmela08,GuGrmelaBous08,hco193}, colloidal suspensions \cite{WagnerN01,GrmelaAitKadiLafl98,Keshtkaretal10,Grmelaetal14}, two-phase systems \cite{Hutter01,Hutter02,Espanol01,EspanolThieu03}), relativistic hydrodynamics \cite{hco109,hco111,hco112,hco114,hco180,Duong15}, discrete formulations of hydrodynamics for simulations \cite{hco122,SerranoEspanol01,EspanolRev03,ElleroEspanolFlek03}, and thermodynamically guided simulations \cite{hco142,hco146,hco145,Kroeger04,hco155,hco157,hco182,IlgKroger11,Ilgetal11,hco195} (see also the review article \cite{hco173} offering ``four lessons and a caveat'' for good simulations in the context of nonequilibrium statistical mechanics).

Several basic transport phenomena have been generalized to the nonlinear regime. For example, diffusion through polymeric and nanocomposite membranes has been modeled by means of the GENERIC framework \cite{LiuDeKee05a,LiuDeKee05b}. Also a comprehensive discussion of the multiscale thermodynamics and mechanics of heat flow goes beyond linear irreversible thermodynamics \cite{GrmelaLebDub11}. Significant progress in applying nonequilibrium thermodynamics to increasingly more complex fluids has, for example, been made with rheological modeling of suspensions of red blood cells \cite{GuGrmelaBous10}, a biological problem of obvious importance. Thermodynamics has also contributed to the understanding of gas flow in the smallest of channels, as in microfluidics, and of aerodynamics of satellites and space stations in the outer limits of our atmosphere \cite{hco194,StruchtrupTorr10,hco198}. Also conceptual developments in electrochemistry benefit from thermodynamic considerations \cite{hco216}.

Whereas the original development of nonequilibrium thermodynamics has mainly been pushed in the context of complex fluids, the general framework is by no means restricted to fluids. Also crystallization phenomena, including polymer crystallization, have been better understood with the help of the methods of modern nonequilibrium thermodynamics and statistical mechanics \cite{Mukherjeeetal04,Hutter04,HutterRutArm05,Hutter06,vanMeerveldetal08,MukherjeeBeris10,BaigEdwards10a,BaigEdwards10b}. Plasticity and viscoplastic solids are further topics in which important issues have been clarified by means of thermodynamics \cite{hco176,HutterTervoort08b,HutterTervoort08,hco192,Mielke11,HuttervBremen12,HutterSvendsen12,HutterTervoort15,Peshkovetal15}. By combining thermodynamics with a thoughtful characterization of the microstructure, valuable insight into continuum damage mechanics has been gained \cite{HutterTervoort08a}. Structural glasses are another challenging problem in physics and materials science for which nonequilibrium phenomena are widely believed to play an important role. Promising new ways to approach this long-standing challenge are suggested or supported by the GENERIC framework \cite{hco167,hco177,hco196,hco208,hco213,FuerederIlg13,SemkivHutter16}.

Most of the applications of nonequilibrium thermodynamics deal with the modeling of bulk systems. To solve the resulting bulk equations one typically needs boundary conditions. The usefulness of linear irreversible thermodynamics for obtaining boundary conditions has been shown by Waldmann in his pioneering 1967 article \cite{Waldmann67}. Brenner and Ganesan \cite{BrennerGane00} asked the very deep question ``Are conditions at a boundary `boundary conditions'?'' Nonequilibrium thermodynamics actually provides the powerful language for expressing the physics at the boundary consistently \cite{hco162,hco171}, thus going well beyond the mathematics of boundary conditions. An illustrative example is provided by the thermodynamic formulation of wall slip \cite{hco172}. Within linear irreversible thermodynamics, a general description of the dynamics of interfaces has been developed by Bedeaux and coworkers \cite{BedeauxAlbMazur76,Bedeaux86,AlbanoBedeaux87}. The main challenge is to generalize the concept of local equilibrium, which is known to be a key ingredient to the nonequilibrium thermodynamics of bulk systems, to lower-dimensional interfaces \cite{JohannessenBed03,JohannessenGroBed08,GlavatskiyBed09,hco204,hco219}. The analysis of the fully nonlinear thermodynamic behavior of complex interfaces within modern nonequilibrium thermodynamics is a very active field of research \cite{hco188,Sagis10,Sagis10a,Sagis11,hco207,Sagis13,Sagis14,SchweizerSagis14,hco223}.

The theory of quantum dissipation \cite{BreuerPetru,Weiss} will play an increasing role in various branches of quantum technology, such as quantum computing, quantum information processing, quantum communication, quantum cryptography, quantum simulation, quantum metrology, quantum sensing, and quantum imaging. After gaining a deep understanding of its geometric structure, modern nonequilibrium thermodynamics suggests the proper form of the equations also for coupling quantum systems to classical environments \cite{Grabert82,hco199,hco221,hco201}. Some applications of the thermodynamic approach to quantum dissipation are discussed in \cite{hco200,hco209,hco217,hco222}. A most remarkable result is that thermodynamics suggests nonlinearity of quantum master equations, contrary to the almost universal presupposition of linear quantum master equations in present-day applications. This nonlinearity can, for example, be handled by stochastic simulation techniques \cite{hco202,hco205}.

\subsection*{A challenge for the future}
According to GENERIC, time-evolution equations for nonequilibrium systems have a well-defined mathematical structure in which reversible and irreversible contributions are identified separately. An important challenge for the future is how the structure of thermodynamically admissible evolution equations can be preserved under time-discretization, which is a key to successful numerical calculations (space-discretization is straightforward because GENERIC can handle discrete systems, as illustrated by the above-mentioned discrete formulations of hydrodynamics for the purpose of simulations). Why should we bother to establish thermodynamically consistent models and boundary conditions if we cannot preserve their features in the numerical solution techniques, which we almost unavoidably have to resort to in most relevant applications? An analogous challenge is well-known to be met by the symplectic integrators used for Hamiltonian systems \cite{SanzSernaCalvo}. Such symplectic integrators are known for better performance in numerical calculations for Hamiltonian systems and hence provide the motivation for similarly introducing GENERIC integrators to perform better calculations for nonequilibrium systems. GENERIC integrators should preserve the time-invariance of the Poisson bracket expressed by the Jacobi identity, the mutual degeneracy requirements, as well as the symmetry and positive-semidefiniteness of the friction matrix. The important advantages of GENERIC integrators hence are:

\noindent (i) The rigorous conservation of an approximate energy leads to limited fluctuations of the physical energy, even over long periods of time.

\noindent (ii) The strict separation of reversible and irreversible contributions to dynamics.

\noindent (iii) A reliable reproduction of the physical entropy production without any spurious dissipation.

Inspired by Moser's work for Hamiltonian systems (see Section~3(e) of \cite{Moser68}), we develop the requirements to be imposed on GENERIC integrators. We then construct GENERIC integrators for systems with a single dissipative process by Taylor expansion. After discussing some details for two specific examples (harmonic oscillator with friction, two gas containers exchanging volume and energy), we summarize the general features of GENERIC integrators and identify some possible directions for future work.

%\bibliography{P:/latex_personal/bibtex/hcopubs}

\end{document}